# Complete Characterizations of Well-Posedness in Parametric Composite Optimization

### v2

@@ Abstract
This paper provides complete characterization of well-posedness for Karush–Kuhn–Tucker (KKT) systems associated with general problems of perturbed composite optimization. Leveraging the property of parabolic regularity for composite models, we show that the second-order subderivative of the cost function reduces to the novel second-order variational function playing a crucial role in the subsequent analysis. This foundational result implies that the strong second-order sufficient condition introduced in this work for the general class of composite optimization problems naturally extends the classical second-order sufficient condition in nonlinear programming. Then we obtain several equivalent characterizations of the second-order qualification condition (SOQC) and highlight its equivalence to the constraint nondegeneracy condition under the $\mathcal{C}^{2}$-cone reducibility assumption. These insights lead us to multiple equivalent conditions for the major Lipschitz-like/Aubin property of KKT systems, including SOQC combined with the new second-order subdifferential condition and SOQC combined with tilt stability of local minimizers. Furthermore, under $\mathcal{C}^{2}$-cone reducibility, we prove that the Lipschitz-like property of the reference KKT system is equivalent to its strong regularity. Finally, we demonstrate that the Lipschitz-like property is equivalent to the nonsingularity of the generalized Jacobian associated with the KKT system under a certain verifiable assumption. These results provide a unified and rigorous framework for analyzing stability and sensitivity of solutions to composite optimization problems, as well as for the design and justification of numerical algorithms.
Keywords: composite optimization, variational analysis and generalized differentiation, well-posedness, parabolic regularity, KKT systems, second-order sufficient conditions, tilt stability
@@ 1 Introduction
This paper concerns the study of the general class of composite optimization problems given by
@@
where $h:\mathcal{R}^{n}\rightarrow\mathcal{R}$, $F:\mathcal{R}^{n}\rightarrow\mathcal{R}^{m}$ are ${\cal C}^{1}$-smooth mappings and $g:\mathcal{R}^{m}\rightarrow\overline{\mathcal{R}}:=\mathcal{R}\cup\{\infty\}$ is an extended-real-valued lower semicontinuous (l.s.c.) convex function, which is always assumed to be proper, i.e., $g(x)>$-$\infty$ for all $x\in\mathcal{R}^{m}$ with ${\rm dom}\,g:=\{x\in\mathcal{R}^{m}\;|\;g(x)<\infty\}\neq\emptyset$. The extended-real-valued format in (1.1) allows us to incorporate various constraints via $F(x)\in{\rm dom}\,g$. In particular, model (1.1) covers problems of nonlinear programming, second-order cone programming, and semidefinite programming that correspond to the indicator function $g$ of the corresponding cone, as well as other classes of constrained problems considered below.
@@
During the past few decades, great many efforts have been made to investigate diverse well-posedness issues for constraint and variational systems under perturbations; see, e.g., the monographs [Bonnans2000, DontchevandRockafellar2014, KlatteKummer2002, Mordukhovich2006, Mordukhovich2018, Mordukhovich2024, Rockafellar1998] and the references therein. In his landmark work [Robinson1980], Robinson introduced the notion of strong regularity for the generalized equation $0\in\Phi(x,u)$ at $(\bar{x},\bar{u})$, which means that the inverse of the linearized mapping
$(x,u)\mapsto\left[\begin{array}[]{c}\nabla_{xx}^{2}L(\bar{x},\bar{u})(x-\bar{x})+\nabla F(\bar{x})^{T}(u-\bar{u})\\ -F(\bar{x})-\nabla F(\bar{x})(x-\bar{x})\end{array}\right]+\left[\begin{array}[]{c}0\\ \partial g^{*}(u)\end{array}\right]$
@@
admits a Lipschitz continuous single-valued localization around the origin, where $A^{T}$” stands for the matrix transposition. Characterizing the strong regularity of perturbed KKT systems at their solution points has emerged as a significant research topic in mathematical programming. In the context of nonlinear programming (NLP), it is established in [Robinson1980, Theorem 4.1] and [BonnansSulem1995, Theorem 4.10] that the strong regularity of the canonically perturbed KKT system is equivalent to the simultaneous fulfillment of the strong second-order sufficient condition (SSOSC) and the linear independence constraint qualification (LICQ), as well as [Robinson1980, Theorem 3.1] to the nonsingularity of the associated Clarke’s generalized Jacobian of the KKT system at the corresponding solution point.
@@
As shown in [DontchevRockafellar1996, Theorem 5], the strong regularity of KKT systems associated with NLPs is equivalent to the Lipschitzian behavior of mappings that was first revealed by Aubin [Aubin1984] in the context of convex programming under the name of “pseudo-Lipschitz property”. This notion, which is a graphical localization of the classical (Hausdorff) Lipschitz continuity of set-valued mappings, is now called the Aubin or Lipschitz-like property to emphasize its Lipschitzian nature (while “pseudo” means “false”). It has been well realized that such Lipschitzian behavior of set-valued mappings/multifunctions is equivalent to the two major well-posedness concepts in nonlinear and variational analysis known as metric regularity and linear openness/covering applied to the mapping inverse; see, e.g., [Mordukhovich2006, Rockafellar1998] for more details and references.
@@
Developing the line of [Robinson1980], the equivalence between the strong regularity and appropriate versions of SSOSC and nondegeneracy conditions is established in [BonnansRamirezC2005, Theorem 30] for second-order cone programming, while the equivalence between the strong regularity and Aubin property for such programs is obtained in [OutrataJivriv2011, Theorem 21] (see also the Errarum in [Opazo]) under the strict complementarity condition and in [ChenChenSunZhu2025, Theorem 4.12] without the latter. For semidefinite programming, similar results are available in [Sun2006, Theorem 4.1] and [ChenChenSunZhang2025, Theorem 4.1]. For general constrained optimization problems, the results of [MordukhovichNghiaRockafellar2015, Theorem 5.6] (see also [Mordukhovich2024, Section 5.5]) establish equivalence among the properties of strong regularity, Lipschitzian strong stability, Lipschitzian full stability, and second-order subdifferential conditions at local minimizers under the reducibility and nondegeneracy conditions. Subsequently, [Mordukhovich2017, Theorem 3.3] derives the equivalence between the notions of Lipschitzian full stability for local minimizers under some nondegeneracy and of strong regularity of the associated KKT system for a class of composite optimization problems involving convex piecewise linear functions. The recent paper [Tangwang2024] proposes a definition of the generalized SSOSC and demonstrates the equivalence between the introduced SSOSC in conjunction with the nondegeneracy condition, the nonsingularity of Clarke’s generalized Jacobian of the nonsmooth system at a given KKT point, and the strong regularity of the KKT point for composite optimization problems. This comprehensive body of research demonstrates fundamental connections between various regularity and stability concepts in optimization theory, providing a unified framework for analyzing well-posedness behavior of optimization problems and their solution structures.
@@
Recall [Bonnans2000] that the nonnegative orthant cone, the second-order cone and the positive semidefinite cone are $\mathcal{C}^{2}$-cone reducible. This naturally raises the question of whether the Lipschitz-like property of the KKT system is equivalent to the strong regularity of solution points for all $\mathcal{C}^{2}$-cone reducible problems and also in more general settings. While significant steps in this direction have been recently made in [ChenChenSunZhu2025, ChenChenSunZhang2025] for second-order cone and semidefinite programming based on the Mordukhovich criterion, it is mentioned in the conclusion of [ChenChenSunZhang2025] that “currently, it is not clear to us how to extend these results to the generic non-polyhedral ${\cal C}^{2}$-cone reducible constrained optimization.” Nevertheless, the latest preprint [sun25] provides the equivalence between the Aubin property and strong regularity for certain ${\cal C}^{2}$-cone reducible generalized equations by using the degree theory, without any appeal to variational analysis.
@@
In this paper, we employ and develop advanced tools of variational analysis and generalized differentiation that allow us to obtain comprehensive characterizations of well-posedness properties of the KKT systems (1.3) with ${\cal C}^{2}$-cone reducible functions $g$. In particular, we establish in this way the equivalence between the strong regularity and Aubin property for the KKT system under the assumptions different from [sun25, Theorem 4] that is applicable to our model when $g$ is the indicator function of a ${\cal C}^{2}$-cone reducible set; see Remark 5.1 for more discussions.
@@
As has been well realized, primal and dual second-order conditions play a pivotal role in characterizing well-posedness of optimization problems. In what follows, we systematically exploit the powerful property of parabolic regularity for the function $g$ in (1.2) that has been largely investigated in the recent publications [MohammadiMordukhovichSarabi2020, Mohammadi2021, Mohammadi2022, Mohammadi2020]. To derive our main results, we provide a precise description of the second-order qualification condition (SOQC) and establish its equivalence to the nondegeneracy condition under the $\mathcal{C}^{2}$-cone reducibility. Leveraging the properties of isolated calmness and tilt stability, we characterize the Aubin property based on three distinct conditions. Under the parabolic regularity of $g$, we show that this property is equivalent to SOQC combined with the new second-order subdifferential condition, and to SOQC combined with the tilt stability of the reference local minimizer for (1.1). Moreover, under the parabolic regularity of $g$ and a certain mild assumption, which holds for many common functions such as the indicator function of the positive semidefinite cone, the nuclear norm function, and the spectral norm functions, we establish the equivalence between the Aubin property and the nonsingularity of the generalized Jacobian associated with the KKT system. These results provide novel and multifaceted perspectives on characterizing Lipschitzian behavior and related well-posedness properties, while opening new avenues for understanding their connections to various stability and regularity conditions in optimization theory and algorithms.
@@
Our notation is standard in variational analysis; see [Mordukhovich2006, Rockafellar1998]. The closed unit ball in $\mathcal{R}^{m}$ is denoted by $\mathbb{B}_{\mathcal{R}^{m}}$. For a given point $x\in\mathcal{R}^{m}$ and a number $\varepsilon>0$, the closed ball centered at $x$ with radius $\varepsilon$ is defined by $\mathbb{B}(x,\varepsilon):=\{u\in\mathcal{R}^{m}\,|\,\|u-x\|\leq\varepsilon\}$. Given a set $C\subset\mathcal{R}^{m}$, the indicator function $\delta_{C}$ of $C$ is defined by $\delta_{C}(x):=0$ if $x\in C$ and $\delta_{C}(x):=\infty$ otherwise. For a point $x\in\mathcal{R}^{m}$, $\operatorname{dist}(x,C)$ denotes the distance from $x$ to the set $C$, while $\operatorname{\Pi}_{C}(x)$ stands for the projection of $x$ onto $C$. For a closed convex cone $C\neq\emptyset$, the lineality space of $C$ is $\operatorname{lin}C:=C\cap(-C)$ representing the largest subspace contained in $C$, and the affine hull $\operatorname{aff}C:=C-C$ is the smallest subspace containing $C$. Given a matrix $A\in\mathcal{R}^{m\times n}$, the kernel of $A$ is $\operatorname{ker}A:=\{x\in\mathcal{R}^{n}\,|\,Ax=0\}$, the range of $A$ is $\operatorname{rge}A$, and $A^{\dagger}$ denotes the Moore-Penrose inverse. The notation $C_{1}\subset C_{2}$ means that $C_{1}$ belongs or equals to $C_{2}$.
@@ 2 Basic Tools of Variational Analysis
In this section, we recall some fundamental tools of variational analysis and generalized differentiation while referring to the monographs [Bonnans2000, Mordukhovich2006, Rockafellar1998] for more details.
@@
When $f$ is convex, it follows from [Rockafellar1998, Proposition 13.37] that (2.3) is single-valued and Lipschitz continuous on $\mathcal{R}^{m}$. The (first-order) subderivative of $f:\mathcal{R}^{m}\rightarrow\overline{\mathcal{R}}$ at $\bar{x}\in{\rm dom}\,f$ in the direction $d\in\mathcal{R}^{m}$ is defined by
@@
It has been realized that the parabolic regularity holds for broad classes of extended-real-valued functions; see [Bonnans2000, MohammadiMordukhovichSarabi2020, Mohammadi2021, Mohammadi2022, Mohammadi2020, Rockafellar1998]. The following proposition demonstrates that the parabolic regularity of a convex function is inherited by the indicator function of its epigraph.
@@
It follows from [Rockafellar1998, Theorem 8.9] that $\lambda(\bar{\mu},-1)\in\partial\delta_{\operatorname{epi}f}(\bar{x},f(\bar{x}))$. Combining [Rockafellar1998, Example 13.62], [Mohammadi2021, Theorem 3.3], [Rockafellar1998, Theorem 8.2], and [Bonnans2000, Proposition 2.126] tells us that
@@
where the last equality is derived from the twice epi-differentiability of $f$ based on [MordukhovichTangWang-tilt-2025, Proposition 3.2]. Moreover, it follows from [Mohammadi2021, Proposition 3.2] and [Bonnans2000, Proposition 3.41] that
$d^{2}\delta_{\operatorname{epi}f}((\bar{x},f(\bar{x})),\lambda(\bar{\mu},-1))(d,\gamma)\leq\inf_{(w,\beta)\in\mathcal{T}^{2}_{\operatorname{epi}f}((\bar{x},f(\bar{x})),(d,\gamma))}\{\lambda\langle\bar{\mu},w\rangle-\lambda\beta\}=\lambda d^{2}f(\bar{x},\bar{\mu})(d).$
@@
$\begin{array}[]{ll}\widehat{\mathcal{N}}_{C}(\bar{x})&:=\big\{d\,\big|\,\langle d,x-\bar{x}\rangle\leq o(\|x-\bar{x}\|)\ \mbox{for}\ x\in C\big\},\\ \mathcal{N}_{C}(\bar{x})&:=\big\{d\,\big|\,\exists\,x^{k}\rightarrow\bar{x}\ \mbox{with}\ x^{k}\in C\ \mbox{and}\ d^{k}\rightarrow d\ \mbox{with}\ d^{k}\in\widehat{\mathcal{N}}_{C}(x^{k})\big\}.\end{array}$ (2.6)
@@
Relationships between the well-posedness properties in Definition 2.2 have been clearly understood in variational analysis; see, e.g., the books [DontchevandRockafellar2014, Mordukhovich2018, Rockafellar1998] and the references therein.
@@
Lemma 2.1 is known as the Mordukhovich criterion; see [Mordukhovich1993, Theorem 5.7] and [Rockafellar1998, Theorem 9.40]. Lemma 2.2 is called the Levy-Rockafellar criterion; see [DontchevandRockafellar2014, Theorem 4E.1] and the references therein. Lemma 2.3 is taken from [Gfrerer2022, Theorem 2.7]. This criterion has been recently enhanced in [BenkoRockafellar2024, Theorem 4.2] in the context where the multifunction $S$ is the solution mapping $S_{\operatorname{KKT}}$ associated with the KKT system in (1.3).

### v3

@@ Abstract
This paper characterizes the well-posedness of Karush-Kuhn-Tucker system for perturbed composite optimization. Using the parabolic regularity, we introduce a novel second-order variational function, shown to be the pivotal object governing second-order behavior. This foundational result yields the strong second-order sufficient condition introduced here for the general class of composite optimization problems to naturally extend the classical second-order sufficient condition in nonlinear programming. Then we obtain several equivalent characterizations of the second-order qualification condition (SOQC) and highlight its equivalence to the constraint nondegeneracy condition under the $\mathcal{C}^{2}$-cone reducibility assumption. These insights lead us to multiple equivalent conditions for the major Lipschitz-like/Aubin property of KKT systems, including SOQC combined with the new second-order subdifferential condition and SOQC combined with tilt stability of local minimizers. Under the $\mathcal{C}^{2}$-cone reducibility, we settle the long-standing question by proving the equivalence between the Aubin property of KKT systems around local minimizers and the classical notion of strong regularity under some additional assumptions. Finally, we demonstrate that the Lipschitz-like property is equivalent to the nonsingularity of the generalized Jacobian associated with the KKT system under a certain verifiable assumption. These results provide a unified and rigorous framework for analyzing stability and sensitivity of solutions to composite optimization problems, as well as for the design and justification of numerical algorithms.
Dedicated to the memory of Franco Giannessi,
an outstanding mathematician and human being
Keywords: variational analysis and generalized differentiation, well-posedness, parabolic regularity, second-order sufficient conditions, tilt stability
@@ 1 Introduction
This paper addresses the areas of variational analysis, generalized differentiation, and stability in optimization in which contributions by Professor Giannessi are difficult to overstate. We particularly concern the study of the general class of composite optimization problems given by
@@
where $h:\mathcal{R}^{n}\rightarrow\mathcal{R}$, $F:\mathcal{R}^{n}\rightarrow\mathcal{R}^{m}$ are ${\cal C}^{2}$-smooth mappings and $g:\mathcal{R}^{m}\rightarrow\overline{\mathcal{R}}:=\mathcal{R}\cup\{\infty\}$ is an extended-real-valued lower semicontinuous (l.s.c.) convex function, which is always assumed to be proper, i.e., $g(x)>$-$\infty$ for all $x\in\mathcal{R}^{m}$ with ${\rm dom}\,g:=\{x\in\mathcal{R}^{m}\;|\;g(x)<\infty\}\neq\emptyset$. The extended-real-valued format in (1.1) allows us to incorporate various constraints via $F(x)\in{\rm dom}\,g$. In particular, model (1.1) covers problems of nonlinear programming, second-order cone programming, and semidefinite programming that correspond to the indicator function $g$ of the corresponding cone, as well as other classes of constrained problems considered below.
@@
During the past few decades, great many efforts have been made to investigate diverse well-posedness issues for constraint and variational systems under perturbations; see, e.g., the monographs [4, 12, 16, 32, 33, 34, 39] and the references therein. In his landmark work [38], Robinson introduced the notion of strong regularity for the generalized equation $0\in\Phi(x,u)$ at $(\bar{x},\bar{u})$, which means that the inverse of the linearized mapping
$(x,u)\mapsto\left[\begin{array}[]{c}\nabla_{xx}^{2}L(\bar{x},\bar{u})(x-\bar{x})+\nabla F(\bar{x})^{T}(u-\bar{u})\\ -F(\bar{x})-\nabla F(\bar{x})(x-\bar{x})\end{array}\right]+\left[\begin{array}[]{c}0\\ \partial g^{*}(u)\end{array}\right]$ (1.5)
@@
admits a Lipschitz continuous single-valued localization around the origin, where $A^{T}$” stands for the matrix transposition. Characterizing the strong regularity of perturbed KKT systems at their solution points has emerged as a significant research topic in mathematical programming. In the context of nonlinear programming (NLP), it is established in [38, Theorem 4.1] and [5, Theorem 4.10] that the strong regularity of the canonically perturbed KKT system is equivalent to the simultaneous fulfillment of the strong second-order sufficient condition (SSOSC) and the linear independence constraint qualification (LICQ), as well as [38, Theorem 3.1] to the nonsingularity of the associated Clarke’s generalized Jacobian of the KKT system at the corresponding solution point.
@@
As shown in [11, Theorem 5], the strong regularity of KKT systems associated with NLPs is equivalent to the Lipschitzian behavior of mappings that was first revealed by Aubin [1] in the context of convex programming under the name of “pseudo-Lipschitz property”. This notion, which is a graphical localization of the classical (Hausdorff) Lipschitz continuity of set-valued mappings, is now called the Aubin or Lipschitz-like property to emphasize its Lipschitzian nature (while “pseudo” means “false”). It has been well realized that such Lipschitzian behavior of set-valued mappings/multifunctions is equivalent to the two major well-posedness concepts in nonlinear and variational analysis known as metric regularity and linear openness/covering applied to the mapping inverse; see, e.g., [32, 39] for more details and references.
@@
Developing the line of [38], the equivalence between the strong regularity and appropriate versions of SSOSC and nondegeneracy conditions is established in [3, Theorem 30] for second-order cone programming, while the equivalence between the strong regularity and Aubin property for such programs is obtained in [36, Theorem 21] (see also the Errarum in [35]) under the strict complementarity condition and in [7, Theorem 4.12] without the latter. For semidefinite programming, similar results are available in [41, Theorem 4.1] and [6, Theorem 4.1]. For general constrained optimization problems, the results of [24, Theorem 5.6] (see also [34, Section 5.5]) establish equivalence among the properties of strong regularity, Lipschitzian strong stability, Lipschitzian full stability, and second-order subdifferential conditions at local minimizers under the reducibility and nondegeneracy conditions. Subsequently, [27, Theorem 3.3] derives the equivalence between the notions of Lipschitzian full stability for local minimizers under some nondegeneracy and of strong regularity of the associated KKT system for a class of composite optimization problems involving convex piecewise linear functions. The recent paper [42] proposes a definition of the generalized SSOSC and demonstrates the equivalence between the introduced SSOSC in conjunction with the nondegeneracy condition, the nonsingularity of Clarke’s generalized Jacobian of the nonsmooth system at a given KKT point, and the strong regularity of the KKT point for composite optimization problems. This comprehensive body of research demonstrates fundamental connections between various regularity and stability concepts in optimization theory, providing a unified framework for analyzing well-posedness behavior of optimization problems and their solution structures.
@@
Recall [4] that the nonnegative orthant cone, the second-order cone and the positive semidefinite cone are $\mathcal{C}^{2}$-cone reducible. This naturally raises the question of whether the Lipschitz-like property of the KKT system is equivalent to the strong regularity of solution points for all $\mathcal{C}^{2}$-cone reducible problems and also in more general settings. While significant steps in this direction have been recently made in [7, 6] for second-order cone and semidefinite programming based on the Mordukhovich criterion, it is mentioned in the conclusion of [6] that “currently, it is not clear to us how to extend these results to the generic non-polyhedral ${\cal C}^{2}$-cone reducible constrained optimization.” Nevertheless, the latest preprint [19] provides the equivalence between the Aubin property and strong regularity for certain ${\cal C}^{2}$-cone reducible generalized equations by using the degree theory, without any appeal to variational analysis.
@@
In this paper, we employ and develop advanced tools of variational analysis and generalized differentiation that allow us to obtain comprehensive characterizations of well-posedness properties of the KKT systems (1.3) with ${\cal C}^{2}$-cone reducible functions $g$. In particular, we establish in this way the equivalence between the strong regularity and Aubin property for the KKT system under certain additional assumptions different from [19, Theorem 4] that is applicable to our model when $g$ is the indicator function of a ${\cal C}^{2}$-cone reducible set.
@@
As has been well realized, primal and dual second-order conditions play a pivotal role in characterizing well-posedness of optimization problems. In what follows, we systematically exploit the powerful property of parabolic regularity for the function $g$ in (1.2) that has been largely investigated in the recent publications [20, 21, 22, 23]. To derive our main results, we provide a precise description of the second-order qualification condition (SOQC) and establish its equivalence to the nondegeneracy condition under the $\mathcal{C}^{2}$-cone reducibility. Leveraging the properties of isolated calmness and tilt stability, we characterize the Aubin property based on three distinct conditions. Under the parabolic regularity of $g$, we show that this property is equivalent to SOQC combined with the new second-order subdifferential condition, and to SOQC combined with the tilt stability of the reference local minimizer for (1.1). Imposing the $\mathcal{C}^{2}$-cone reducibility condition, we prove that the Aubin property of the reference KKT system, being combined with the assumption that the primal part of the local solution to the canonically perturbed KKT is a local minimizer of (1.1), is equivalent to its strong regularity. Moreover, under the parabolic regularity of $g$ and a certain mild assumption, which holds for many common functions such as the indicator function of the positive semidefinite cone, the nuclear norm function, and the spectral norm functions, we establish the equivalence between the Aubin property and the nonsingularity of the generalized Jacobian associated with the KKT system. These results provide novel and multifaceted perspectives on characterizing Lipschitzian behavior and related well-posedness properties, while opening new avenues for understanding their connections to various stability and regularity conditions in optimization theory and algorithms.
@@
Our notation is standard in variational analysis; see [32, 39]. The closed unit ball in $\mathcal{R}^{m}$ is denoted by $\mathbb{B}_{\mathcal{R}^{m}}$. For a given point $x\in\mathcal{R}^{m}$ and a number $\varepsilon>0$, the closed ball centered at $x$ with radius $\varepsilon$ is defined by $\mathbb{B}(x,\varepsilon):=\{u\in\mathcal{R}^{m}\,|\,\|u-x\|\leq\varepsilon\}$. Given a set $C\subset\mathcal{R}^{m}$, the indicator function $\delta_{C}$ of $C$ is defined by $\delta_{C}(x):=0$ if $x\in C$ and $\delta_{C}(x):=\infty$ otherwise. For a point $x\in\mathcal{R}^{m}$, $\operatorname{dist}(x,C)$ denotes the distance from $x$ to the set $C$, while $\operatorname{\Pi}_{C}(x)$ stands for the projection of $x$ onto $C$. For a closed convex cone $C\neq\emptyset$, the lineality space of $C$ is $\operatorname{lin}C:=C\cap(-C)$ representing the largest subspace contained in $C$, and the affine hull $\operatorname{aff}C:=C-C$ is the smallest subspace containing $C$. Given a matrix $A\in\mathcal{R}^{m\times n}$, the kernel of $A$ is $\operatorname{ker}A:=\{x\in\mathcal{R}^{n}\,|\,Ax=0\}$, the range of $A$ is $\operatorname{rge}A$, and $A^{\dagger}$ denotes the Moore-Penrose inverse. The notation $C_{1}\subset C_{2}$ means that $C_{1}$ belongs or equals to $C_{2}$.
@@ 2 Basic Tools of Variational Analysis
In this section, we recall some fundamental tools of variational analysis and generalized differentiation while referring to the monographs [4, 32, 39] for more details.
@@
When $f$ is convex, it follows from [39, Proposition 13.37] that (2.3) is single-valued and Lipschitz continuous on $\mathcal{R}^{m}$. The (first-order) subderivative of $f:\mathcal{R}^{m}\rightarrow\overline{\mathcal{R}}$ at $\bar{x}\in{\rm dom}\,f$ in the direction $d\in\mathcal{R}^{m}$ is defined by
@@
It has been realized that the parabolic regularity holds for broad classes of extended-real-valued functions; see [4, 20, 21, 22, 23, 39]. The following proposition demonstrates that the parabolic regularity of a convex function is inherited by the indicator function of its epigraph.
@@
It follows from [39, Theorem 8.9] that $\lambda(\bar{\mu},-1)\in\partial\delta_{\operatorname{epi}f}(\bar{x},f(\bar{x}))$. Combining [39, Example 13.62], [21, Theorem 3.3], [39, Theorem 8.2], and [4, Proposition 2.126] tells us that
@@
where the last equality is derived from the twice epi-differentiability of $f$ based on [29, Proposition 3.2]. Moreover, it follows from [21, Proposition 3.2] and [4, Proposition 3.41] that
$d^{2}\delta_{\operatorname{epi}f}((\bar{x},f(\bar{x})),\lambda(\bar{\mu},-1))(d,\gamma)$ $\leq\inf_{(w,\beta)\in\mathcal{R}^{m}\times\mathcal{R}}\left\{d^{2}\delta_{\operatorname{epi}f}(\bar{x},f(\bar{x}))((d,\gamma),(w,\beta))-\lambda\langle\bar{\mu},w\rangle+\lambda\beta\right\}$
$=\inf_{(w,\beta)\in\mathcal{T}^{2}_{\operatorname{epi}f}((\bar{x},f(\bar{x})),(d,\gamma))}\{\lambda\langle\bar{\mu},w\rangle-\lambda\beta\}=\lambda d^{2}f(\bar{x},\bar{\mu})(d).$
@@
$\begin{array}[]{ll}\widehat{\mathcal{N}}_{C}(\bar{x}):=\big\{d\,\big|\,\langle d,x-\bar{x}\rangle\leq o(\|x-\bar{x}\|)\ \mbox{for}\ x\in C\big\},\\ \mathcal{N}_{C}(\bar{x}):=\big\{d\,\big|\,\exists\,x^{k}\rightarrow\bar{x}\ \mbox{with}\ x^{k}\in C\ \mbox{and}\ d^{k}\rightarrow d\ \mbox{with}\ d^{k}\in\widehat{\mathcal{N}}_{C}(x^{k})\big\}.\end{array}$ (2.6)
@@
Relationships between the well-posedness properties in Definition 2.2 have been clearly understood in variational analysis; see, e.g., the books [12, 33, 39] and the references therein.
@@
Lemma 2.1 is known as the Mordukhovich criterion; see [31, Theorem 5.7] and [39, Theorem 9.40]. Lemma 2.2 is called the Levy-Rockafellar criterion; see [12, Theorem 4E.1] and the references therein. Lemma 2.3 is taken from [13, Theorem 2.7]. This criterion has been recently enhanced in [2, Theorem 4.2] in the context where the multifunction $S$ is the solution mapping $S_{\operatorname{KKT}}$ associated with the KKT system in (1.3).